# Multi-Review Fusion-in-Context

### v1

@@ E.4 Additional Meta Evaluation Results
Table 8: Statistics of our dataset. The two right-most columns indicate the percentage of summary sentences that align with spans in more than one review, and the percentage of summary sentences that align with a non-continuous span from across more than one review sentence in one of its reviews.
@@ Appendix H Additional Experimental Details
We also set the maximum target length to 200, which we found works best, as well as setting the batch size to 1. The other parameters are similar to Slobodkin et al. (2022, 2023a). The model is trained for 10k steps. Training is performed on a two A100-SXM4-80GB GPUs, and costs about 12 GPU hours for the supervised models (Flan-T5H, Flan-T5no-H, and Flan-T5only-H) and about 36 GPU hours for the RL-tuned variant of Flan-T5H.
@@ Appendix I Generation Examples
Table 9 illustrates the performance on the test set of flan-t5 with different model sizes, finetuned on the FiC trainswt.

### v2

@@ E.4 Additional Meta Evaluation Results
Table 8: Full statistics of our dataset, including the number of unique review-sets, the average number of summaries per review-set, the number of summary/review-set pairs (a unique review-set creates a pair with each of its summaries), the mean review/summary size (in tokens and in sentences), the maximum review/review-set/summary size (in tokens), the percentage of summary sentences whose alignments span across more than one review, and the percentage of summary sentences whose alignments span across more than one review sentence within one of its reviews (namely, within a single review, the alignments come from more than one sentence).
@@ Appendix H Additional Experimental Details
We also set the maximum target length to 200, which we found works best, as well as setting the batch size to 1. The other parameters are similar to Slobodkin et al. (2022, 2023a). The model is trained for 10k steps. Training is performed on two A100-SXM4-80GB GPUs, and costs about 12 GPU hours for the supervised models (Flan-T5H, Flan-T5no-H, and Flan-T5only-H) and about 36 GPU hours for the RL-tuned variant of Flan-T5H.
@@ Appendix I Generation Examples
Table 9 illustrates the performance on the test set of flan-t5 with different model sizes, finetuned on the FiC trainset.